# Quality and Trust in LLM-generated Code

### 09 Feb 2024 v2

@@ 3.2.2. Line-level Code Completion
Code completion is a well-studied task in Software Engineering, and widely deployed in IDEs, with tools like GitHub Copilot (Chen et al., 2021). This task can be investigated at both the token and line levels (Izadi et al., 2022; Kim et al., 2021; Lu et al., 2021). Performance on this task provides a good indication of model’s ability to generate a sequence of code tokens, conditioned on a particular context (viz., “auto-regressive” generation); usually the preceding sequence of prompt tokens. In real development settings, developers may not always require the complete function from the model. Instead, generating the next several tokens is often helpful enough. Full-function code synthesis is more challenging than small-scale completion; moreover, partial code completion is already in wide use, and has been evaluated using real software development data (Ziegler et al., 2022; Tabachnyk and Nikolov, 2022); however, calibration for code-completion, which is the most widely-deployed task for LLMs, has so far not been evaluated in detail. The current decoder-only GPT models (Brown et al., 2020; Chen et al., 2021) are already trained to generate the next token at low average cross-entropy given all the prior tokens, following the condition $p(token\ |\ prior\ tokens)$.
Unsurprisingly, we found that the model is per se well-calibrated at the token level. In this work, we will primarily focus on line-level completion. While several datasets exist for this problem (Allamanis and Sutton, 2013; Raychev et al., 2016), we use DyPyBench (Bajaj, 2023b, a), a new dataset, consisting of 50 popular open-source Python projects, including test suites for these projects. The test suites allow a test-correctness measure, in addition to the highly restrictive exact-match (viz., the original line). Figure 3b presents a sample prompt and model output for the line completion task, where we provide the model with all the prior lines and ask it to generate the next line.
DyPyBench consists of complex real-world projects, each with hundreds of thousands of lines of Python code, and totaling over 2.2 million lines of Python code [^7] As per naive sum of lines in all non-virtualenv *.py files across all project directories. We ran all test suites for each project with coverage reporting enabled. We subsequently extracted all functions from the projects using our derivative of function_parser introduced by (Husain et al., 2019), explored the dataset to ensure a lack of duplicates, and then joined each function with its line-level coverage report data. To make evaluation feasible, we randomly selected 1,988 functions with a body line span greater than 3, $100\%$ test coverage, and at least one line in the “Docstring”.
@@ 3.2.3. Program Repair
Program repair is a well-studied problem in Software Engineering (Goues et al., 2019). Several studies have applied LLMs to the program repair problem and shown that it is effective (Fan et al., 2023; Jiang et al., 2023; Ahmed and Devanbu, 2023). However, LLM calibration for program repair is not well-understood. Several program-repair datasets are available, including the widely-used Defects4J (Just et al., 2014), which includes real-world examples of buggy programs, with fixes and test-sets. This paper primarily focuses on line-level bug repair, represented-by 120 samples from the Defects4J dataset. However, with only 120 samples, we may not obtain a comprehensive view of calibration. Therefore, we included another dataset, ManySStubs4J (Karampatsis and Sutton, 2020) (abbr. SStubs), which consists of single-line repairs. Considering resource constraints, we selected 3,000 samples from this dataset, which is sufficient to investigate calibration. However, ManySStubs4J does not provide test-sets; so the only evaluation metric available is the exact-matching of the generated text to the ground truth bug-free text.
@@ 3.3.1. Codex
We used three models for our experiments; OpenAI Code-davinci-002, which we refer to as Codex, is one of them. Codex is a family of GPT-3 based high-performance code models (Chen et al., 2021), which is used in GitHub Copilot products. The Codex is trained on large volumes of both natural language and source code from publicly available sources, including GitHub. Codex is most capable in Python, but it is proficient in over a dozen languages, including JavaScript, Go, Perl, PHP, Ruby, Swift, TypeScript, and even Shell [^9] https://openai.com/blog/openai-codex. We use Code-davinci-002 in our experiments, as it is the most capable version of the Codex family, with 175 billion parameters trained on a more recent text and code corpus compared to previous members of the family. It is a completion model; we access it using an API key. OpenAI provides logits for each of the generated tokens, from which we calculate token probabilities (See  § 3.1.2).
@@ 3.3.3. CodeGen2
We also studied an open-source model called CodeGen2 (Nijkamp et al., 2023). CodeGen2 is an enhanced version of the CodeGen (Nijkamp et al., 2022) model, and its authors have aimed to improve code synthesis performance by unifying four key components: (i) model architectures, (ii) learning methods, (iii) infill sampling, and (iv) data distributions. CodeGen2 is available in various sizes (1B, 3.7B, 7B, and 16B parameters), but we specifically utilized the largest and best performing 16-billion-parameter variant in our experiments. However, even the largest CodeGen2 model is much smaller than the OpenAI models we examine in this paper.
@@ 3.4. Summary of Complete Pipeline
    Solving: LLM Code Generation: The LLM is invoked to generate a solution for the coding prompt: Python code for DyPyBench, HumanEval, and MBPP datasets, and Java for the Defects4J and SStubs datasets. This step either uses the OpenAI API, or (in the case of CodeGen2), a local A100 80GB GPU. Post-solution, confidence measures are computed for the generated code which are used in our analysis. For patching & line completion tasks, all reference solutions are $<80$ tokens (median $\approx 10$ tokens). For function synthesis tasks, all references are $<440$ tokens (median of $\approx 40$ tokens, p90 of $\approx 115$ tokens). Therefore, we choose to sample up to a maximum of 500 tokens. We sample from the model with temperature=0, consistent with the reality that busy developers typically look at just the first suggestion in the completion task  (Barke et al., 2022). For function synthesis task, temperature zero is most accurate and is fairly standard practice when doing pass@1 with only one solution to generate and present  (Chen et al., 2021; Liu et al., 2023). In the case of function synthesis problems, we stop generation at the end of the prompted function block. In the case of line-level problems, we stop at a newline character. The completion is then merged into its original context. For example, in DyPyBench, the generated line replaces only the original line in the file, the project’s test suite is run in the DyPyBench Docker environment, ultimately producing a test report which is parsed and recorded.
@@ 4.1.1. Overall Correctness
It is interesting to note that approximately half the test-passing generations did not match the original code exactly; furthermore 6.89% of the cases where the code matched exactly, did not pass all the test cases. Upon careful study we found that these tests were “flaky”, depending on network conditions, execution order, and other variable execution environment conditions. This aligns with (Bajaj, 2023a), the author of this dataset, who observed an overall $8.07\%$ failure rate, but noted that 31 out of 50 projects had zero failed tests. This table illustrates the relative merits/demerits of each correctness-evaluating approach, in practical SE settings. Since the correctness performance is different with these two notions of correctness, the calibration is also different, as we see below.
@@ 4.1.3. Confidence Measures
The intrinsic vs reflective measure values have no consistent relationship, even for a given model, dataset and task. This lack of relationship may not necessarily be negative: e.g., perhaps the model’s reflective, prompted confidence may be better calibrated, as suggested by prior work (Bai et al., 2022; Austin et al., 2021).
@@ 4.2.3. Is Rescaling a Panacea for Calibration?
As discussed in § 4.2.2, there is a clear trend that rescaling does improve calibration (RQ5). It has been used in settings other than generative models of code, with other notions of correctness  (Guo et al., 2017; Minderer et al., 2021; Desai and Durrett, 2020; Park and Caragea, 2022; Chen et al., 2023a; Liang et al., 2023; Li et al., 2019). However, there are disadvantages. First, (see  § 4.2.2), “bucket collapse” can mislead with deceptively low ECE. Second, some correctness data is needed, to fit rescaling parameters.
@@ 4.2.5. RQ3: How well do reflective calibration approaches work?
Alternative and improved ways of prompting (e.g., few-shotting (Kadavath et al., 2022a), different verbalization formats (Lin et al., 2022; Tian et al., 2023), fine-tuning (Kadavath et al., 2022a; Lin et al., 2022), chain-of-thought (Wei et al., 2022; Tian et al., 2023), etc.) may alter these findings and are areas for future work.
@@
We discuss this question with respect to DyPyBench, since it is a realistic, large-sample dataset supporting both notions of correctness. As discussed earlier (§ 4.1.2), OpenAI models perform better overall (regarding correctness) on test-passing than on exact-match; but they are better calibrated for exact-match! Here, the OpenAI models’ intrinsic $p_{tot}$ probability exhibits very high skill for predicting exact-match correctness, in the range of 0.31 to 0.43 (appendix Table 7). The same Codex $p_{tot}$ shows the highest skill for test-passing correctness, but with much lower Skill Scores (around 0.09). Thus the measure can skillfully predict when it repeats the exact sequence of human-written code found in the corpus, but it is less skilled in predicting when the code is correctly test-passing. This kind of “data recovery” could be expected, given the original training objective of LLMs (Carlini et al., 2020). It is also noteworthy that the exact-match skill for the same models for the SStubs dataset is lower (max ~ 0.06), possibly because these are cases where the human programmer originally made mistakes, and as reported by Jesse et al. (2023), Codex tends to prefer to repeat those errors, rather than their fixes. Thus the exact-match correctness in SStubs is more demanding, because the desired code in this bug-fix dataset is different from the originally human-written versions.
@@ 5. Discussion
Language models are now widely-integrated into Software Engineering practice, via tools like Copilot (Ziegler et al., 2022) and Didact (Large sequence models for software development activities, [n. d.]). We raise here the importance of calibration when integrating generative LLMs in coding practice. We evaluate the calibration of generative LLM use (especially code completion) with large samples of realistic data (DyPyBench, SStubs), using widely adopted models, as well as some more academic datasets (HumanEval, MBPP).
@@
We’ve always had bugs; poor-quality code isn’t new. Our push for calibration, however, arises from the increasing amount of code generated by LLM. GitHub claims that up to 61% of code [^11] https://github.blog/2023-02-14-github-copilot-now-has-a-better-ai-model-and-new-capabilities/ in some systems is generated by LLMs. It is also known that LLMs make a lot of mistakes. A recent paper has reported (Jesse et al., 2023) that LLMs, even when trained on properly fixed code, tends to recapitulate the the old unfixed, buggy code when prompted in context. However, LLMs do have very high capacity, and a demonstrated ability to usefully reflect (Shinn et al., 2023; Bai et al., 2022) on their generated text. Thus we have both a high risk (of buggy code), and a chance to improve productivity. We believe that improved calibration could lead to better management of the risk-benefit of LLM-generated code. The studied correctness calibration is a stepping stone for more complex notions of confidence (like severity and localized confidence). Additionally, by studying code LLMs, we might make progress on the general safe deployment of capable generative models (Gros et al., 2023).
@@ 6. Threats to Validity
Our exploration is not exhaustive; other SE tasks and datasets also could benefit from calibration studies. Additionally, the specific prompts we used for this paper surely played a role in our findings. Other prompts or problem phrasings (such as different forms of context for line-level code completion) may yield different results. Regarding test flakiness: the test “flake” rate of DyPyBench is not zero, but is not unrealistic (Luo et al., 2014). Additionally, conclusions may shift as models continue to scale and improve, with possibly emergent capabilities.
@@ 7. Related Work
LLMs for code are extensively studied (Zhang et al., 2023a; Zheng et al., 2023). While calibration has a long history in modeling (Brier, 1950; Steyerberg et al., 2010), it is not a frequently studied topic in the SE community. Early work moving into modern machine learning studied the calibration of smaller neural models performing classification tasks on text and images; while these early models were poorly calibrated per se, their performance could be improved by simple scaling (Guo et al., 2017) of their output probabilities. As models became larger, calibration was found to improve (Srivastava et al., 2022). Pre-training was also found to improve calibration (Hendrycks et al., 2019; Desai and Durrett, 2020); however, these findings have been disputed (Chen et al., 2023b).
@@
More recent works evaluated LLM calibration on a wide variety of settings (Kadavath et al., 2022b; Jiang et al., 2021; Desai and Durrett, 2020; Key et al., 2022).  Desai and Durrett (2020) studied non-code (natural language) tasks such as inference or paraphrasing, with only intrinsic measures using older-generation models (BERT and RoBERTA). Jiang et al. (2021) studied calibration for natural language question-answering using just intrinsic measures. In contrast, we study calibration for three coding-related tasks, using both artificial and natural code datasets, and both intrinsic and reflective confidence measures, to evaluate calibration in the SE domain.
@@
Li et al. (2019) investigate the calibration of Computer vision (CV) models from an operational perspective i.e., the shift between training input and production inputs, presenting it as a software quality problem that can be addressed using Bayesian approaches. Minderer et al. (2021) evaluate the calibration of at the time, state of the art CV models and find improved calibration with more recent models, notably those not using convolutions. Park and Caragea (2022) study the effect of the mixup technique (Zhang et al., 2018) on calibration in a natural language understanding (NLU) setting using older generation models (BERT and RoBERTa). Chen et al. (2023a) investigate the calibration of pretrained language models on various NLP tasks, also using older generation models (RoBERTa and T5). Liang et al. (2023) introduce the HELM benchmark, which includes calibration as one of its seven metrics to evaluate language models in a natural language context.
@@
When suitably prompted, Kadavath et al. (2022b) found that LLMs can output well-calibrated scores on whether their own answers are correct or not, viz., larger models “know what they know”. While this work did investigate some function synthesis tasks (HumanEval & an unpublished Python function dataset), they did so using only their private models, and ultimately focused on natural language tasks. Key et al. (2022) developed an approach that given a natural language problem description, produces a confidence score for a sampled candidate solution based on generated specifications, allowing them to judge whether the LLM can solve the problem at all. Recent work has also explored calibration of software topics such as root cause analysis (Zhang et al., 2023c).

### 16 Feb 2024 v3

@@ 3.2.2. Line-level Code Completion
Code completion is a well-studied task in Software Engineering, and widely deployed in IDEs, with tools like GitHub Copilot (chen2021evaluating, 12). This task can be investigated at both the token and line levels (izadi2022codefill, 23, 33, 41). Performance on this task provides a good indication of model’s ability to generate a sequence of code tokens, conditioned on a particular context (viz., “auto-regressive” generation); usually the preceding sequence of prompt tokens. In real development settings, developers may not always require the complete function from the model. Instead, generating the next several tokens is often helpful enough. Full-function code synthesis is more challenging than small-scale completion; moreover, partial code completion is already in wide use, and has been evaluated using real software development data (ziegler2022productivity, 71, 57); however, calibration for code-completion, which is the most widely-deployed task for LLMs, has so far not been evaluated in detail. The current decoder-only GPT models (brown2020language, 10, 12) are already trained to generate the next token at low average cross-entropy given all the prior tokens, following the condition $p(token\ |\ prior\ tokens)$.
Unsurprisingly, we found that the model is per se well-calibrated at the token level. In this work, we will primarily focus on line-level completion. While several datasets exist for this problem (allamanis2013mining, 2, 51), we use DyPyBench (bajaj_piyush_2023_7886366, 6, 7), a new dataset, consisting of 50 popular open-source Python projects, including test suites for these projects. The test suites allow a test-correctness measure, in addition to the highly restrictive exact-match (viz., the original line). Figure 3b presents a sample prompt and model output for the line completion task, where we provide the model with all the prior lines and ask it to generate the next line.
DyPyBench consists of complex real-world projects, each with hundreds of thousands of lines of Python code, and totaling over 2.2 million lines of Python code [^7] As per naive sum of lines in all non-virtualenv *.py files across all project directories. We ran all test suites for each project with coverage reporting enabled. We subsequently extracted all functions from the projects using our derivative of function_parser introduced by (codesearchnet, 22), explored the dataset to ensure a lack of duplicates, and then joined each function with its line-level coverage report data. To make evaluation feasible, we randomly selected 1,988 functions with a body line span greater than 3, $100\%$ test coverage, and at least one line in the “Docstring”.
@@ 3.2.3. Program Repair
Program repair is a well-studied problem in Software Engineering (goues2019automated, 18). Several studies have applied LLMs to the program repair problem and shown that it is effective (fan2023automated, 17, 26, 1). However, LLM calibration for program repair is not well-understood. Several program-repair datasets are available, including the widely-used Defects4J (just2014defects4j, 28), which includes real-world examples of buggy programs, with fixes and test-sets. This paper primarily focuses on line-level bug repair, represented-by 120 samples from the Defects4J dataset. However, with only 120 samples, we may not obtain a comprehensive view of calibration. Therefore, we included another dataset, ManySStubs4J (karampatsis2020often, 31) (abbr. SStubs), which consists of single-line repairs. Considering resource constraints, we selected 3,000 samples from this dataset, which is sufficient to investigate calibration. However, ManySStubs4J does not provide test-sets; so the only evaluation metric available is the exact-matching of the generated text to the ground truth bug-free text.
@@ 3.3.1. Codex
We used three models for our experiments; OpenAI Code-davinci-002, which we refer to as Codex, is one of them. Codex is a family of GPT-3 based high-performance code models (chen2021evaluating, 12), which is used in GitHub Copilot products. The Codex is trained on large volumes of both natural language and source code from publicly available sources, including GitHub. Codex is most capable in Python, but it is proficient in over a dozen languages, including JavaScript, Go, Perl, PHP, Ruby, Swift, TypeScript, and even Shell [^9] https://openai.com/blog/openai-codex. We use Code-davinci-002 in our experiments, as it is the most capable version of the Codex family, with 175 billion parameters trained on a more recent text and code corpus compared to previous members of the family. It is a completion model; we access it using an API key. OpenAI provides logits for each of the generated tokens, from which we calculate token probabilities (See  § 3.1.2).
@@ 3.3.3. CodeGen2
We also studied an open-source model called CodeGen2 (nijkamp2023codegen2, 45). CodeGen2 is an enhanced version of the CodeGen (nijkamp2022codegen, 46) model, and its authors have aimed to improve code synthesis performance by unifying four key components: (i) model architectures, (ii) learning methods, (iii) infill sampling, and (iv) data distributions. CodeGen2 is available in various sizes (1B, 3.7B, 7B, and 16B parameters), but we specifically utilized the largest and best performing 16-billion-parameter variant in our experiments. However, even the largest CodeGen2 model is much smaller than the OpenAI models we examine in this paper.
@@ 3.4. Summary of Complete Pipeline
    Solving: LLM Code Generation: The LLM is invoked to generate a solution for the coding prompt: Python code for DyPyBench, HumanEval, and MBPP datasets, and Java for the Defects4J and SStubs datasets. This step either uses the OpenAI API, or (in the case of CodeGen2), a local A100 80GB GPU. Post-solution, confidence measures are computed for the generated code which are used in our analysis. For patching & line completion tasks, all reference solutions are $<80$ tokens (median $\approx 10$ tokens). For function synthesis tasks, all references are $<440$ tokens (median of $\approx 40$ tokens, p90 of $\approx 115$ tokens). Therefore, we choose to sample up to a maximum of 500 tokens. We sample from the model with temperature=0, consistent with the reality that busy developers typically look at just the first suggestion in the completion task  (Barke_James_Polikarpova_2022, 8). For function synthesis task, temperature zero is most accurate and is fairly standard practice when doing pass@1 with only one solution to generate and present  (chen2021evaluating, 12, 39). In the case of function synthesis problems, we stop generation at the end of the prompted function block. In the case of line-level problems, we stop at a newline character. The completion is then merged into its original context. For example, in DyPyBench, the generated line replaces only the original line in the file, the project’s test suite is run in the DyPyBench Docker environment, ultimately producing a test report which is parsed and recorded.
@@ 4.1.1. Overall Correctness
It is interesting to note that approximately half the test-passing generations did not match the original code exactly; furthermore 6.89% of the cases where the code matched exactly, did not pass all the test cases. Upon careful study we found that these tests were “flaky”, depending on network conditions, execution order, and other variable execution environment conditions. This aligns with (bajaj2023, 7), the author of this dataset, who observed an overall $8.07\%$ failure rate, but noted that 31 out of 50 projects had zero failed tests. This table illustrates the relative merits/demerits of each correctness-evaluating approach, in practical SE settings. Since the correctness performance is different with these two notions of correctness, the calibration is also different, as we see below.
@@ 4.1.3. Confidence Measures
The intrinsic vs reflective measure values have no consistent relationship, even for a given model, dataset and task. This lack of relationship may not necessarily be negative: e.g., perhaps the model’s reflective, prompted confidence may be better calibrated, as suggested by prior work (bai2022constitutional, 5, 4).
@@ 4.2.3. Is Rescaling a Panacea for Calibration?
As discussed in § 4.2.2, there is a clear trend that rescaling does improve calibration (RQ5). It has been used in settings other than generative models of code, with other notions of correctness  (guo2017calibration, 20, 43, 16, 49, 13, 37, 36). However, there are disadvantages. First, (see  § 4.2.2), “bucket collapse” can mislead with deceptively low ECE. Second, some correctness data is needed, to fit rescaling parameters.
@@ 4.2.5. RQ3: How well do reflective calibration approaches work?
Alternative and improved ways of prompting (e.g., few-shotting (Kadavath2022LanguageM, 29), different verbalization formats (lin2022teaching, 38, 58), fine-tuning (Kadavath2022LanguageM, 29, 38), chain-of-thought (Wei2022ChainOT, 61, 58), etc.) may alter these findings and are areas for future work.
@@
We discuss this question with respect to DyPyBench, since it is a realistic, large-sample dataset supporting both notions of correctness. As discussed earlier (§ 4.1.2), OpenAI models perform better overall (regarding correctness) on test-passing than on exact-match; but they are better calibrated for exact-match! Here, the OpenAI models’ intrinsic $p_{tot}$ probability exhibits very high skill for predicting exact-match correctness, in the range of 0.31 to 0.43 (appendix Table 7). The same Codex $p_{tot}$ shows the highest skill for test-passing correctness, but with much lower Skill Scores (around 0.09). Thus the measure can skillfully predict when it repeats the exact sequence of human-written code found in the corpus, but it is less skilled in predicting when the code is correctly test-passing. This kind of “data recovery” could be expected, given the original training objective of LLMs (Carlini2020ExtractingTD, 11). It is also noteworthy that the exact-match skill for the same models for the SStubs dataset is lower (max ~ 0.06), possibly because these are cases where the human programmer originally made mistakes, and as reported by (10174227, 25, ), Codex tends to prefer to repeat those errors, rather than their fixes. Thus the exact-match correctness in SStubs is more demanding, because the desired code in this bug-fix dataset is different from the originally human-written versions.
@@ 5. Discussion
Language models are now widely-integrated into Software Engineering practice, via tools like Copilot (ziegler2022productivity, 71) and Didact (researchLargeSequence, 35). We raise here the importance of calibration when integrating generative LLMs in coding practice. We evaluate the calibration of generative LLM use (especially code completion) with large samples of realistic data (DyPyBench, SStubs), using widely adopted models, as well as some more academic datasets (HumanEval, MBPP).
@@
We’ve always had bugs; poor-quality code isn’t new. Our push for calibration, however, arises from the increasing amount of code generated by LLM. GitHub claims that up to 61% of code [^11] https://github.blog/2023-02-14-github-copilot-now-has-a-better-ai-model-and-new-capabilities/ in some systems is generated by LLMs. It is also known that LLMs make a lot of mistakes. A recent paper has reported (10174227, 25) that LLMs, even when trained on properly fixed code, tends to recapitulate the the old unfixed, buggy code when prompted in context. However, LLMs do have very high capacity, and a demonstrated ability to usefully reflect (shinn2023reflexion, 54, 5) on their generated text. Thus we have both a high risk (of buggy code), and a chance to improve productivity. We believe that improved calibration could lead to better management of the risk-benefit of LLM-generated code. The studied correctness calibration is a stepping stone for more complex notions of confidence (like severity and localized confidence). Additionally, by studying code LLMs, we might make progress on the general safe deployment of capable generative models (gros2023ai, 19).
@@ 6. Threats to Validity
Our exploration is not exhaustive; other SE tasks and datasets also could benefit from calibration studies. Additionally, the specific prompts we used for this paper surely played a role in our findings. Other prompts or problem phrasings (such as different forms of context for line-level code completion) may yield different results. Regarding test flakiness: the test “flake” rate of DyPyBench is not zero, but is not unrealistic (luo2014empirical, 42). Additionally, conclusions may shift as models continue to scale and improve, with possibly emergent capabilities.
@@ 7. Related Work
LLMs for code are extensively studied (Zhang2023ASO, 67, 69). While calibration has a long history in modeling (brier1950verification, 9, 56), it is not a frequently studied topic in the SE community. Early work moving into modern machine learning studied the calibration of smaller neural models performing classification tasks on text and images; while these early models were poorly calibrated per se, their performance could be improved by simple scaling (guo2017calibration, 20) of their output probabilities. As models became larger, calibration was found to improve (srivastava2022beyond, 55). Pre-training was also found to improve calibration (hendrycks2019using, 21, 16); however, these findings have been disputed (chen-etal-2023-close, 14).
@@
More recent works evaluated LLM calibration on a wide variety of settings (kadavath2022language, 30, 27, 16, 32).  (desai2020calibration, 16, ) studied non-code (natural language) tasks such as inference or paraphrasing, with only intrinsic measures using older-generation models (BERT and RoBERTA). (jiang2021can, 27, ) studied calibration for natural language question-answering using just intrinsic measures. In contrast, we study calibration for three coding-related tasks, using both artificial and natural code datasets, and both intrinsic and reflective confidence measures, to evaluate calibration in the SE domain.
@@
(Li2019OperationalCD, 36, ) investigate the calibration of Computer vision (CV) models from an operational perspective i.e., the shift between training input and production inputs, presenting it as a software quality problem that can be addressed using Bayesian approaches. (Minderer2021RevisitingTC, 43, ) evaluate the calibration of at the time, state of the art CV models and find improved calibration with more recent models, notably those not using convolutions. (Park2022OnTC, 49, ) study the effect of the mixup technique (Zhang_Cisse_Dauphin_Lopez, 66) on calibration in a natural language understanding (NLU) setting using older generation models (BERT and RoBERTa). (Chen_Yuan_Cui_Liu_Ji_2023, 13, ) investigate the calibration of pretrained language models on various NLP tasks, also using older generation models (RoBERTa and T5). (Liang2023HolisticEO, 37, ) introduce the HELM benchmark, which includes calibration as one of its seven metrics to evaluate language models in a natural language context.
@@
When suitably prompted, (kadavath2022language, 30, ) found that LLMs can output well-calibrated scores on whether their own answers are correct or not, viz., larger models “know what they know”. While this work did investigate some function synthesis tasks (HumanEval & an unpublished Python function dataset), they did so using only their private models, and ultimately focused on natural language tasks. (Key_Li_Ellis_2022, 32, ) developed an approach that given a natural language problem description, produces a confidence score for a sampled candidate solution based on generated specifications, allowing them to judge whether the LLM can solve the problem at all. Their metrics include calibration. Recent work has also explored calibration of software topics such as root cause analysis (zhang2023pace, 65).